OpenStudy (anonymous):

Hello Math people! please help me!? I will give best Response and A Medal! thank you! Evaluate x + y when x = –94 and y = –24. (Points : 1) –118 –70 70 118

OpenStudy (anonymous):

first choice

OpenStudy (anonymous):

x+y=-94-24=-118

OpenStudy (anonymous):



OpenStudy (studygurl14):

-118

OpenStudy (anonymous):

Yes!!! Thank you so Much!!!